OpenStudy (anonymous):

solve equation by graphing, give an answer with 2 decimal... 10^2-4-3x

OpenStudy (anonymous):

you mean 10x^2?

OpenStudy (anonymous):

sorry its 10x^20=4-3x

OpenStudy (anonymous):

oh that is a tough question, the equation cannot be solved exactly

OpenStudy (anonymous):

why is that ? would 4x^2-x=6 ???

OpenStudy (anonymous):

you cannot solve a equation with a x^20, that terrible :S

OpenStudy (anonymous):

i must have been confused maybe it is 10x^2=4-3x

OpenStudy (anonymous):

ok...its x=-4/5 or x=1/2